How many atoms of oxygen are in 350 mg of calcium nitrate?

2.07 × 1026

9.48 × 1025

7.7 × 1021

2.1 × 1026

1.3 × 1021
Answer #1

mass of calcium nitrate = 350 mg = 0.350 g

number of moles of calcium nitrate = 0.350 g / 164.088 g/mol = 0.00213 mole

1 mole of calcium nitrate = 6.023*10^23 atoms so

0.00213 mole of calcium nitrate =

= 0.00213 mole of calcium nitrate *(6.023*10^23 atoms / 1 mole of calcium nitrate)

= 1.28*10^21 atoms

Therefore, the number of atoms of calcium nitrate = 1.28*10^21